# On the Sharp Input-Output Analysis of Nonlinear Systems under Adversarial Attacks

### v1

@@ 2 Problem Formulation
$w_{t}$ is an adversarial attack at each time $t$ with probability $p<\frac{1}{2\tau}$ conditioned on $\mathcal{F}_{t}$. In other words, the adversary uses the information in the filtration $\mathcal{F}_{t}$ to generate disturbances $w_{t}$, which can therefore be correlated and possibly adversarial. The attack probability specifies that the system is not under attack with probability $1-p$, meaning that
$\mathbb{P}(w_{t}\equiv 0~{}|~{}\mathcal{F}_{t})=1-p$ (10)
holds for all $t=0,\dots,T-1$ and $\mathcal{F}_{t}=\bm{\sigma}\{x_{0},u_{0},\dots,u_{t},w_{0},\dots,w_{t-1}\}$.
@@
Our goal is to accurately estimate the true matrix $G^{*}$ that describes the input-output mapping from the control inputs $(u_{t},\dots,u_{t-\tau})$ to the observation $y_{t}$. Recall that the relationship (6) and (7) can be written as
$y_{t}=G^{*}\cdot\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}$ (11)
for all $t\geq\tau$. Given a time horizon $T$, we aim to learn the true system $G^{*}$ using the following $\ell_{1}$-norm estimator based on partial observations $y_{\tau},\dots,y_{T-1}$ and control inputs $u_{0},\dots,u_{T-1}$:
$\hat{G}_{T}=\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|y_{t}-G\cdot\Phi(\bm{U}_{t}^{(\tau)})\right\|_{1}$ (12)
Under the stated assumptions, we will show in the next section that the $\ell_{1}$-norm estimator achieves the optimal estimation error $\Theta(\rho^{\tau})$, where $\rho$ is the contraction factor in Assumption 2.1 and $\tau$ is the input memory length.
@@ 3.1 Main Theorem
Suppose that Assumptions 2.1, 2.3, 2.4, 2.6, and 2.8 hold. Consider $\nu:=\frac{\sqrt{M}\tau L_{\phi}\sigma_{u}}{\lambda}$. Let $G^{*}$ be the true matrix governing the system (11) and $\hat{G}_{T}$ denote a solution to the $\ell_{1}$-norm estimator given in (12). Given $\delta\in(0,1]$, when
$T\geq\Theta\left(\frac{\tau^{2}\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\left[M\log\left(\frac{\nu}{2(1-p)^{\tau}-1}\right)+\log\left(\frac{r}{\delta}\right)\right]\right),$ (13)
@@
$\|G^{*}-\hat{G}_{T}\|_{F}\leq\Theta\left(\frac{\rho^{\tau}L\sqrt{r}}{\lambda}\cdot\frac{(\sigma_{u}+\sigma_{w})\tau\nu^{3}}{(1-\rho)(2(1-p)^{\tau}-1)}\right)$ (14)
@@
Our main theorem states that after the time given in (13), the estimation error of $\Theta(\rho^{\tau})$ is achieved, considering that additional polynomial terms in $\tau$ are dominated by the exponential decay in $\tau$. However, notice that the estimation error does not decay as the time $T$ increases, and thus cannot converge to zero. While this error bound decreases as $\tau$ grows, the memory length $\tau$ will be chosen as a finite number at the user’s discretion, and thus the bound should be treated as a positive constant. This suggests that the user may want to choose a sufficiently long $\tau$ to obtain a smaller estimation error. However, increasing $\tau$ has three drawbacks: First, it restricts the attack probability as stated in Remark 2.9. Second, the required time (13) implies that it takes longer to arrive at the desired estimation estimation bound. Also, the basis function $\Phi$ may become significantly complex to incorporate longer history, and naturally the optimization problem needs far more computations. Thus, even though the estimation error may decrease with increasing $\tau$, the aforementioned demerits create an inherent trade-off in selecting an appropriate value for $\tau$.
@@ 3.2 Analysis Outline
3.2.1 Analysis Without Past State Effect
Our proof technique starts from a special case where the term $\bm{x}_{t}^{(\tau)}$ in the equation (11) is zero. This auxiliary setting will later be generalized to the case where $\bm{x}_{t}^{(\tau)}$ is nonzero. In the following theorem, we establish a sufficient condition for the true matrix $G^{*}$ to be the unique solution to the $\ell_{1}$-norm minimization problem (12).
@@ 3.2.1 Analysis Without Past State Effect
Let $\bm{W}_{t}^{(\tau,i)}$ denote the $i^{th}$ coordinate of $\bm{W}_{t}^{(\tau)}$ for $i=1,\dots,r$. Suppose that $\bm{x}_{t}^{(\tau)}=0$ for all $t$. Then, $G^{*}$ is the unique solution to the $\ell_{1}$-norm estimator (12) if
$\sum_{t=\tau}^{T-1}\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau,i)}=0\}+s^{T}\Phi(\bm{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})>0,\quad\forall s\in\mathbb{S}^{M-1}$ (15)
holds for all $i\in\{1,\dots,r\}$.
@@
Theorem 3.3 implies that if the left-hand side given in equation (15) is positive for all $s\in\mathbb{S}^{M-1}$ and all coordinates $i$, we will actually be able to exactly recover the true matrix $G^{*}$ with the $\ell_{1}$-norm estimator. As a first step, we prove that for a fixed $s$, the left-hand side of (15) will be sufficiently positive within a finite time. In particular, when $T\geq 2\tau$, we show that
$\sum_{t=\tau}^{T-1}\bigr{|}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\bigr{|}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau,i)}=0\}\geq\Theta(T)$ (16)
with high probability in Lemma D.2 under our assumptions on control inputs: sub-Gaussianity (Assumption 2.3) and Lipschitzness and excitement of the system (Assumption 2.6). We now consider a scenario in which $\mathbb{P}(\bm{W}_{t}^{(\tau,i)}>0)=\mathbb{P}(\bm{W}_{t}^{(\tau,i)}<0)$. Under this sign-symmetry assumption along with the same assumptions on control inputs, Lemma D.4 specifies that
$\left\|\sum_{t=\tau}^{T-1}s^{T}\Phi(\bm{U}_{t}^{(\tau)})\cdot\text{sgn}(\bm{W}_{t}^{(\tau,i)})\right\|_{\psi_{2}}\leq\Theta(\sqrt{T}),$ (17)
where one can use Hoeffding’s inequality (see (29) in Appendix B) to ensure that the sum of the terms in (16) and (17) is sufficiently positive in the order of $\Theta(T)$ with high probability for a fixed $s$.
@@
We now analyze how the term in (15) changes when evaluated at two different points $s,\tilde{s}\in\mathbb{S}^{M-1}$. We show that the difference is lower-bounded by $-\Theta(T\|s-\tilde{s}\|_{2})$ in Lemma D.6, meaning that the closer the points, the smaller the discrepancy. Thus, if one can select a sufficient number of points in $\mathbb{S}^{M-1}$ that make the term in (15) simultaneously positive with high probability, then their surrounding neighborhoods will also yield positive values, implying that the term in (15) is positive for all points in $\mathbb{S}^{M-1}$. Lemma D.7 specifies how many points are needed to achieve universal positivity, which is attained within finite time as shown in Lemma D.8.
Note that the above analysis was made under the sign symmetry $\mathbb{P}(\bm{W}_{t}^{(\tau,i)}>0)=\mathbb{P}(\bm{W}_{t}^{(\tau,i)}<0)$. To relax this assumption to incorporate any adversarial disturbances, we impose the condition $\mathbb{P}(\bm{W}_{t}^{(\tau)}\equiv 0)>0.5$. Then, our analysis is extended to the asymmetric case, in the sense that any distribution that satisfies $\mathbb{P}(\bm{W}_{t}^{(\tau)}\equiv 0)>0.5$ is the approximation of a sign-symmetric distribution where the magnitudes of a certain subset of the support are arbitrarily small, as shown in Lemma D.9. Corollary D.10 states that under the attack probability $p<\frac{1}{2\tau}$, we have $\mathbb{P}(\bm{W}_{t}^{(\tau)}\equiv 0)\geq(1-p)^{\tau}>0.5$. This implies that the term in (15) is bounded below by a positive constant in the order of $\Theta(T)$, and as a by-product, the estimation error is exactly zero within finite time due to Theorem 3.3. ∎
3.2.2 Beyond the Zero Past State Effect
@@
In general, the term $\bm{x}_{t}^{(\tau)}$ in (11) is nonzero, as $\|x_{t-\tau}\|_{2}$ is not zero (see Lemma 2.5). Thus, we need to extend the previous analysis in Section 3.2.1 to general cases. From the optimality of $\hat{G}_{T}$ for the $\ell_{1}$-norm estimator (12) and the input-output mapping (11), we have
$\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}\|_{1}\leq\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}\|_{1},$ (18)
@@
$\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{1}-\|\bm{W}_{t}^{(\tau)}\|_{1}\leq 2\sum_{t=\tau}^{T-1}\|\bm{x}_{t}^{(\tau)}\|_{1},$ (19)
where the left-hand side turns out to be the perturbation of $\ell_{1}$-norm estimator without the past state effect $\bm{x}_{t}^{(\tau)}$. This can be lower-bounded by using the positive constant lower bound $\Theta(T)$ of the term in (15) constructed in the previous Section 3.2.1. The right-hand side is also upper-bounded by $\Theta(T)$ due to Lemma D.11. Accordingly, we can bound the estimation error $\|G^{*}-\hat{G}_{T}\|_{F}$ using (19) and obtain the results in Theorem 3.1 (see the detailed proof in Theorem D.12). ∎
@@ 3.3 Lower Bound
Given $\delta\in(0,1]$, suppose that adversarial attacks $w_{t}$ are designed by an adversary to satisfy $\sigma_{w}\geq\bigr{(}\frac{1}{\rho}\bigr{)}^{\Theta(\tau\log(T/\delta))}$, where $10<\rho<$ 1 is the contraction factor of and $\tau$ is the input memory length. Then, there exists a problem instance satisfying Assumptions 2.1, 2.3, 2.4, 2.6, and 2.8 that suffers from $\Theta\bigr{(}\frac{\rho^{\tau}L\sqrt{r}}{\lambda}\bigr{)}$ estimation error with probability at least $1-\delta$ for any estimator.
Proof Sketch. Under the attack probability $\frac{1}{2\tau+1}$, which satisfies Assumption 2.8, the maximum consecutive attack-free length is bounded by $\Theta(\tau\log(T/\delta))$ with probability at least $1-\delta$. Let the initial state $x_{0}$ be $\sigma_{w}$. Then, the adversarial attacks with $\sigma_{w}$ enable the property $x_{t}\geq 1$ for all $t$ with high probability. This implies that $y_{t}$ in (2) can be written in two different functions $h_{1}\neq h_{2}$ such that
@@
where $x_{t-\tau}$ and the disturbances are set to 00. Choose the functions $h_{1}$ and $h_{2}$ to have different function values for (21), while satisfying the equation (20) for all $x_{t-\tau}\geq 1$. As a result, any estimator may recover either one of the mappings $h_{1}$ or $h_{2}$ arbitrarily, given the same observation trajectory $y_{0},y_{1},\dots,y_{T-1}$. In particular, the two expressions in (21) can differ by $\Theta(\rho^{\tau}L)$, leading to an estimation error $\Theta(\frac{\rho^{\tau}L}{\lambda})$. Finally, this scalar system can be naturally extended to the $r$-dimensional system to yield the stated lower bound. The proof details can be found in Appendix E. ∎
@@
We have established the lower bound $\Theta(\frac{\rho^{\tau}L\sqrt{r}}{\lambda})$, which implies that the estimation error is bounded below by a positive constant for any estimator due to adversarial attacks. While this matches the upper bound (14) up to the same order, the assumption on the sub-Gaussian norm of the attacks depends on $T$. If this norm is required to be uniformly bounded over all $T\geq 0$, it remains unclear whether the gap between the upper and lower bounds can be further tightened. Meanwhile, in Appendix E, we have developed nonlinear basis functions to achieve the desired lower bound. It remains an intriguing open question whether there exists a linear system instance for which the upper and lower bounds match under the constraint that all basis functions are linear in control inputs.

### v2

@@ 2 Problem Formulation
$w_{t}$ is an adversarial attack at each time $t$ with probability $p<\frac{1}{2\tau}$ conditioned on $\mathcal{F}_{t}$; i.e., there exists a sequence $(\xi_{t})_{t\geq 0}$ of independent $\text{Bernoulli}(p)$ variables, each independent of any $\mathcal{F}_{t}$, such that
$\{\xi_{t}=0\}\subseteq\{w_{t}=0\},\quad\forall t\geq 0.$ (11)
Assumption 2.8 specifies that the system is not under attack (case 1 in Figure 1(b)) with probability at least 1p1-p, since $\xi_{t}=0$ implies $w_{t}=0$. At attack times (case 2 in Figure 1(b)), the adversary uses the information in the filtration $\mathcal{F}_{t}$ to generate disturbances $w_{t}$, which can therefore be correlated and possibly adversarial.
@@
Given a time horizon $T$, we aim to learn the true system $G^{*}$ in (8), using the following $\ell_{2}$-norm estimator based on partial observations $y_{\tau},\dots,y_{T-1}$ and control inputs $u_{0},\dots,u_{T-1}$:
$\hat{G}_{T}=\operatorname*{arg\,min}_{G}\sum_{t=\tau}^{T-1}\left\|y_{t}-G\cdot\Phi(\bm{U}_{t}^{(\tau)})\right\|_{2}$ (12)
Under the stated assumptions, we will show in the next section that the $\ell_{2}$-norm estimator achieves the optimal estimation error $O(\rho^{\tau})$, where $\rho$ is the contraction factor in Assumption 2.1 and $\tau$ is the input memory length.
@@ 3.1 Main Theorem
Suppose that Assumptions 2.1, 2.4, 2.5, 2.6, and 2.8 hold, and that the approximation error vector satisfies $\|\bm{\epsilon}_{t}\|_{2}\leq\bar{\epsilon}$ for all $t$. Consider $\nu:=\frac{\sqrt{M\tau}L_{\phi}\sigma_{u}}{\lambda}$. Let $G^{*}$ be the true matrix governing the system (8) and $\hat{G}_{T}$ denote a solution to the $\ell_{2}$-norm estimator given in (12). Given $\delta\in(0,1]$, when
$T=\Omega\left(\frac{\tau\nu^{8}}{(2(1-p)^{\tau}-1)^{2}}\left[rM\log\left(\frac{\tau\nu}{2(1-p)^{\tau}-1}\right)+\log\left(\frac{1}{\delta}\right)\right]\right),$ (13)
@@
$\|G^{*}-\hat{G}_{T}\|_{F}=O\left(\left(\frac{\rho^{\tau}L}{\lambda}\cdot\frac{\sigma_{u}+\sigma_{w}}{1-\rho}+\frac{\bar{\epsilon}}{\lambda}\right)\cdot\frac{\nu^{3}}{2(1-p)^{\tau}-1}\right)$ (14)
@@
Our main theorem states that after the time given in (13) and with a sufficiently small $\bar{\epsilon}$, the estimation error of $O(\rho^{\tau})$ is achieved, considering that additional polynomial terms in $\tau$ are dominated by the exponential decay in $\tau$. However, notice that the estimation error does not decay as the time $T$ increases, and thus cannot converge to zero. While this error bound decreases as $\tau$ grows, the memory length $\tau$ will be chosen as a finite number at the user’s discretion, and thus the bound should be treated as a positive constant. This suggests that the user may want to choose a sufficiently long $\tau$ to obtain a smaller estimation error. However, increasing $\tau$ has three drawbacks: First, it restricts the attack probability as stated in Remark 2.9. Second, the required time (13) implies that it takes longer to arrive at the desired estimation estimation bound. Third, the basis function $\Phi$ may become significantly complex to incorporate longer history, and naturally the optimization problem needs far more computations. Thus, even though the estimation error may decrease with increasing $\tau$, the aforementioned demerits create an inherent trade-off in selecting an appropriate value for $\tau$.
@@ 3.2 Analysis Outline
3.2.1 Analysis Without Past State and Approximation Effect
Our proof technique starts from a special case where the term $\bm{x}_{t}^{(\tau)}$ and $\bm{\epsilon}_{t}$ in the equation (8) are zero. This auxiliary setting will later be generalized to the case where they can take nonzero values. In the following theorem, we establish a sufficient condition for the true matrix $G^{*}$ to be the unique solution to the $\ell_{2}$-norm minimization problem (12).
@@ 3.2.1 Analysis Without Past State and Approximation Effect
Suppose that $\bm{x}_{t}^{(\tau)}=0$ and $\bm{\epsilon}_{t}=0$ for all $t$. Then, $G^{*}$ is the unique solution to the $\ell_{2}$-norm estimator (12) if
$\sum_{t=\tau}^{T-1}\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau)}=0\}-\sum_{t=\tau}^{T-1}\bigr\|Z\Phi(\bm{U}_{t}^{(\tau)})\bigr\|_{2}\cdot\mathbb{I}\{\bm{W}_{t}^{(\tau)}\neq 0\}>0,$ (15)
holds for all $Z\in\mathbb{R}^{r\times M}$ such that $\|Z\|_{F}=1$.
@@
Theorem 3.3 implies that if the left-hand side given in equation (15) is positive for all $Z\in\mathbb{R}^{r\times M}$ such that $\|Z\|_{F}=1$, we will actually be able to exactly recover the true matrix $G^{*}$ with the $\ell_{2}$-norm estimator. In particular, thanks to Lemma 2.3 and (11), we have
$\mathbb{P}(\bm{W}_{t}^{(\tau)}=0)\geq\mathbb{P}(w_{t-1}=0,\dots,w_{t-\tau}=0)\geq\mathbb{P}(\xi_{t-1}=0,\dots,\xi_{t-\tau}=0)=(1-p)^{\tau}>0.5.$
Then, for a fixed $Z$, the sub-Gaussianity of control inputs $u_{t}$, Lipschitzness of $\Phi(\cdot)$, and the excitation condition (10) ensures that the left-hand side of (15) will be sufficiently positive after a finite time.
@@
We now analyze how the term in (15) changes when evaluated at two different points $Z,\tilde{Z}\in\mathbb{R}^{r\times M}$. We show that the difference is indeed small when the points are close. Thus, if one can select a sufficient number of points for which the term in (15) is simultaneously positive with high probability, then their surrounding neighborhoods will also yield positive values. This implies that the term in (15) is universally positive for all points in $\mathbb{R}^{r\times M}$ with unit Frobenius norm. To quantify how many such points are needed, we invoke a well-known covering number argument (Vershynin, 2018). ∎
3.2.2 Beyond the Zero Past State and Approximation Effect
@@
In general, $\bm{x}_{t}^{(\tau)}$ in (8) is nonzero since $\|x_{t-\tau}\|_{2}\neq 0$ (see Lemma 2.3). Moreover, we inevitably face a nonzero approximation error vector $\bm{\epsilon}_{t}$, whose magnitude depends on the expressiveness of the chosen basis functions. Thus, we need to extend the previous analysis in Section 3.2.1 to general cases. From the optimality of $\hat{G}_{T}$ for the $\ell_{2}$-norm estimator (12) and the input-output mapping (8), we have
$\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}+\bm{\epsilon}_{t}\|_{2}\leq\sum_{t=\tau}^{T-1}\|\bm{W}_{t}^{(\tau)}+\bm{x}_{t}^{(\tau)}+\bm{\epsilon}_{t}\|_{2},$ (16)
@@
$\sum_{t=\tau}^{T-1}\|(G^{*}-\hat{G}_{T})\Phi(\bm{U}_{t}^{(\tau)})+\bm{W}_{t}^{(\tau)}\|_{2}-\|\bm{W}_{t}^{(\tau)}\|_{2}\leq 2\sum_{t=\tau}^{T-1}\left(\|\bm{x}_{t}^{(\tau)}\|_{2}+\|\bm{\epsilon}_{t}\|_{2}\right)$ (17)
where the left-hand side turns out to be the perturbation of $\ell_{2}$-norm estimator without the effect of $\bm{x}_{t}^{(\tau)}$ and $\bm{\epsilon}_{t}$. This can be lower-bounded by using the positive constant lower bound $\Omega(T)$ of the term in (15) constructed in the previous Section 3.2.1. The right-hand side is also upper-bounded by $O(T)$ due to Lemma D.9. Accordingly, we can bound the estimation error $\|G^{*}-\hat{G}_{T}\|_{F}$ using (17) to obtain the results in Theorem 3.1. ∎
Remark 3.4.
We note that any $\ell_{\alpha}$-norm estimator with $\alpha\geq 1$ can ensure the left-hand side of (15) remains universally positive even though $\ell_{2}$-norm is replaced by other norms. However, the resulting estimation error bound is weaker than that of the $\ell_{2}$-norm estimator. We analyze two different cases:
[…]
Consequently, the $\ell_{2}$-norm estimator yields a tighter estimation bound $O\bigr(\frac{\rho^{\tau}L}{\lambda}\bigr)$ than those based on other norms since it does not depend on the observation dimension $r$. This will further be supported by the lower bound presented in the next subsection, which is also independent of $r$.
@@ 3.3 Lower Bound
Given $\delta\in(0,1]$, suppose that adversarial attacks $w_{t}$ are designed by an adversary to satisfy $\sigma_{w}=\bigr(\frac{1}{\rho}\bigr)^{\Omega(\tau\log(T/\delta))}$, where $0<\rho<1$ is the contraction factor of $f$, and $\tau$ is the input memory length. Then, there exists a problem instance satisfying Assumptions 2.1, 2.4, 2.5, 2.6, and 2.8 that suffers from $\Omega\bigr(\frac{\rho^{\tau}L}{\lambda}\bigr)$ estimation error with probability at least $1-\delta$ for any estimator.
Proof Sketch. Consider the case where the approximation error vector $\bm{\epsilon}_{t}=0$. Under the attack probability $O(1/\tau)$ in Assumption 2.8, the maximum consecutive attack-free length is bounded by $O(\tau\log(T/\delta))$ with probability at least $1-\delta$. Then, the adversarial attacks enable the property $x_{t}\geq 1$ for all $t$ with high probability. This implies that $y_{t}$ in (2) can be written in two different functions $h_{1}\neq h_{2}$ such that
@@
where $x_{t-\tau}$ and the disturbances are set to 0. Choose the functions $h_{1}$ and $h_{2}$ to have different function values for (19), while satisfying the equation (18) for all $x_{t-\tau}\geq 1$. As a result, any estimator may recover either one of the mappings $h_{1}$ or $h_{2}$ arbitrarily, given the same observation trajectory $y_{0},y_{1},\dots,y_{T-1}$. In particular, the two expressions in (19) can differ by $\Omega(\rho^{\tau}L)$, leading to an estimation error $\Omega(\frac{\rho^{\tau}L}{\lambda})$. The proof details can be found in Appendix E. ∎
@@
We have established the lower bound $\Omega(\frac{\rho^{\tau}L}{\lambda})$, which implies that the estimation error is bounded below by a positive constant for any estimator due to adversarial attacks. While this matches the upper bound (14) up to the same order, the assumption on the sub-Gaussian norm of the attacks depends on $T$. If this norm is required to be uniformly bounded over all $T\geq 0$, it remains unclear whether the gap between the upper and lower bounds can be further tightened. Meanwhile, the proof in Appendix E relies on specially designed nonlinear basis functions to achieve the desired lower bound. It remains an intriguing open question whether there exists a linear system instance for which the upper and lower bounds match under the constraint that all basis functions are linear in control inputs.